# Super-Resolution works for coastal simulations

### v1

@@ III-B Proposed DNCSR for coastal super-resolution
In Figure 2, we show the overall structure of our proposed DNCSR . Given the input coarse simulation data $\mathbf{X}\in\mathbb{R}^{T\times H\times W\times C}$, where T is the number of adjacent views, H and W are the height and width, and $C=3$ is the $U$, $V$, $\xi$ planes. The proposed DNCSR learns the mapping function $f\colon\mathbf{X}\rightarrow\mathbf{Y}$, where $\mathbf{Y}\in\mathbb{R}^{\alpha T\times H\times W\times C}$ is the high-resolution coastal estimation, and $\alpha$ is the temporal upsampling factor. Note that we do not change the spatial resolution because the low-resolution and high-resolution coastal images are rendered from the graph with the same scale but with coarse and fine mesh resolutions, respectively. That is, the resolution of the simulation video does not change, but it becomes more detailed. Inside the network are three key components: Feature extraction, Spatiotemporal attention, and Feature split and reconstruction. Let us introduce them in detail.
@@
Figure 2: The overall structure of the proposed DNCSR . We show the complete architecture of our proposed DNCSR. Given two consecutive coastal simulations, we take them as input to first extract the deep feature representation via multiple RCAB blocks. Then we use spatiotemporal attention to learn the pixel correlations across space and time. Finally, we split the features into three channels for temporal interpolation and spatial super-resolution.
III-C Feature extraction
@@ III-D Spatiotemporal attention
The key component of the DNCSR is the spatiotemporal attention. Our goal is to jointly super-resolve the coarse coastal images spatially and temporally. PDEs typically govern spatiotemporal dynamical systems. Let us first consider a simplified wave equation as
@@
III-E Feature split and reconstruction
In the feature split and reconstruction process, we first project the learned features into the frequency domain using cosine and sine operators, then combine these components to enhance the original features. We subsequently apply a simple convolution layer to extend these features and split them into three subsets, each representing one of the three temporal feature maps for the coastal data. The input consists of two low-resolution coastal images, with the proposed network (DNCSR ) tasked with predicting the intermediate image. To achieve this, we use bilinear interpolation for temporal frame interpolation on the initial feature maps. A short connection is implemented to ensure that the spatiotemporal attention mechanism effectively learns the residuals. Finally, a shared convolution operation is employed to directly output the predicted super-resolution coastal images. This approach ensures that the model captures both spatial and temporal details, producing high-quality intermediate frames.
III-F Losses for coastal super-resolution
To train the proposed DNCSR , we propose utilizing three loss terms to better constrain the visual consistency in space and time, including MAE (Mean Absolute Errors) loss, $L_{p}$ loss, and physics-informed loss. Given the ground truth and estimated super-resolution results ${\mathbf{Y},\mathbf{Y^{\prime}}}\in\mathbb{R}^{\alpha T\times H\times W\times C}$, we first have the MAE loss as,
@@ III-F Losses for coastal super-resolution
[…]
@@
IV Experiments
IV-A Implementation Details
$\bullet$ Datasets. To generate sufficient data samples for model training and analysis, we use UTBEST [10] to simulate the tidal circulation in two locations: Bahamas (Bight of Abaco) and the Galveston Bay. The geometry of both computational domains, bathymetry (bottom topography), and the coarsest computational meshes are illustrated in Figure 4. Finer meshes were obtained by subdividing each triangle into four via edge bisection.
@@ IV-A Implementation Details
At the open sea boundary, both test problems use the tidal elevation prescribed by the function
@@
$+0.1\cos\left(\frac{t}{12.66}+5.93\right)+0.395\cos\left(\frac{t}{12.42}\right)$
$+0.06\cos\left(\frac{t}{12.00}+0.75\right)$
with time $t$ in hours and utilize a quadratic bottom friction $\tau_{bf}=C_{f}|\mathbf{q}|/H^{2}$ with $C_{f}=0.009$ (Bahamas) and $C_{f}=0.004$ (Galveston). The Coriolis parameter for Bahamas and Galveston is set to $3.19\times 10^{-5}~{}\mathrm{s}^{-1}$ and $7.07\times 10^{-5}~{}\mathrm{s}^{-1}$, respectively (see the sources in the introduction for a full description of the test problems).
The simulations are run for 24 hours with the time step for the lowest approximation order (piecewise constant DG spaces) and the lowest mesh resolution set to 20s for the Bahamas and 12s for the Galveston dataset. Output files with the water height $\xi$ and depth-integrated horizontal velocities $U$ and $V$ are generated in each time step for all configurations. Increasing the mesh resolution or the polynomial discretization order requires smaller time steps (we halve the time step for each mesh refinement and each order increment) and results in more output files. Table I summarizes the data used for training with the number of mesh elements shown in ’Resolution’ and the number of time steps in ’Order’ column. We use three approximation orders and resolutions to produce corresponding coarse and fine simulation data, that is, given three approximation orders and three different resolutions, we have a total of $3\times 3$ simulation cases for each location. Then, we render all data of their three views ($U$, $V$, $\xi$ planes) into grayscale images. Finally, all three views are merged as RGB images as the training data.
TABLE I: Summary of the data used in coastal simulation for Bahamas and Galveston. The ’Resolution’ value represents the number of mesh elements. In the ’Order’ column we list the number of time steps in the 24 hour simulation for the corresponding polynomial discretization order.
@@
We randomly split the data into training, validation, and testing in $6:2:2$ ratios without overlapping. The rendered images are of identical dimensions, approximately $900\times 500$ pixels. We also mask out the land area and only focus on the water-covered subdomain super-resolution. In the training phase, we randomly select two adjacent LR coastal images ($\mathbf{X}(t_{n}),\mathbf{X}(t_{n+1})$) to predict three corresponding HR images $\mathbf{Y}(t_{2n}),\mathbf{Y}(t_{2n+1}),\mathbf{Y}(t_{2n+2})$ and crop the image pairs as patches of size $64\times 64$. We randomly apply horizontal & vertical flips, rotations, and frame reversal for data augmentation.
@@
$\bullet$ Parameter setting. We train DNCSR using Adam optimizer with the learning rate of $1\times 10^{-4}$. The learning rate is halved after 30k iterations. The batch size is set to 24, and DNCSR is trained for 100k iterations (about 16 hours) on a PC with one NVIDIA V100 GPU using the PyTorch deep learning platform.
$\bullet$ Metrics and evaluation. We use several metrics for evaluation: MSE (Mean Squared Errors) and MAE (Mean Absolute Errors) measure the average pixel differences between ground truth and estimation. SSIM [74] measures the structural similarity between ground truth and estimation. GMSD [75] measures the gradient differences for perceptual quality assessment. LPIPS [76] measures the perceptual similarity between ground truth and estimation. To measure the model complexity, we use FLOPs, memory, and the number of parameters to compare the computation efficiency.
Figure 5: Speed and performance comparison. We show the overall comparison among different methods in terms of speed (runtime on the vertical axis), performance (RMSE on the horizontal axis), and model complexity (size of the bubbles).
TABLE II: Comparison with state-of-the-art methods on model complexity. We report the results on the Bahamas dataset under the same hardware settings.
Method Number of parameters (M) FLOPs (GMac) Runtime (s)
ST-Bicubic - - 0.003
VDSR* 0.651 8.23 0.038
RDN* 21.99 90.14 0.48
ResShift 118.59 32156 15.2
SwinIR 11.72 1878.41 4.2
EDVR* 20.6 3020.1 1.10
VRT 35.5 3305.75 2.4
PhySR* 0.817 9.20 0.23
Ours 8.07 65.15 1.02
IV-B Comparison of the spatiotemporal super-resolution with state-of-the-art methods
@@
To show the efficiency of our proposed method, we compare it with four state-of-the-art image super-resolution methods: VDSR [77], RDN [63], StableSR [34] and SwinIR [33], two video super-resolution methods: EDVR [52] and VRT [53], and one physics-informed super-resolution method: PhySR [64]. Note that image super-resolution cannot conduct frame interpolation, so we apply spatiotemporal Bicubic to interpolate the missing frames and then apply frame-by-frame super-resolution. All approaches are reimplemented using the publicly available codes, and the methods marked with * indicate that they were retrained from our coastal dataset for a fair comparison. In Table III, both Galveston and Bahamas have two super-resolution schemes. The ’Resolution’ column lists the number of grid elements. The result for each scheme is the average of all three polynomial approximation orders. We can see that our approach achieves the best performance with respect to the RMSE, MAE, SSIM, and GMSD metrics and is the second-best in LPIPS.
@@
TABLE III: Comparison with state-of-the-art methods on coastal super-resolution. We report the deviations from the ground truth given by the highest accuracy simulations for the test datasets of Bahamas and Galveston using different metrics.
Dataset Bahamas Galveston
Resolution 1696 $\rightarrow$ 6784 6784 $\rightarrow$ 27136 3250 $\rightarrow$ 13000 13000 $\rightarrow$ 54000
@@
ST-Bicubic 0.077 0.035 0.964 0.0413 0.108 0.040 0.0208 0.979 0.019 0.097 0.039 0.002 0.972 0.043 0.108 0.620 0.016 0.979 0.079 0.079
VDSR* [77] 0.079 0.033 0.960 0.045 0.118 0.004 0.018 0.975 0.039 0.105 0.020 0.012 0.990 0.023 0.046 0.050 0.016 0.981 0.059 0.048
RDN* [63] 0.054 0.023 0.971 0.046 0.125 0.021 0.009 0.987 0.026 0.105 0.021 0.007 0.989 0.036 0.033 0.030 0.010 0.984 0.049 0.043
ResShift [78] 0.066 0.041 0.977 0.046 0.133 0.023 0.010 0.988 0.025 0.112 0.023 0.012 0.989 0.037 0.040 0.032 0.014 0.986 0.047 0.048
SwinIR [33] 0.045 0.026 0.975 0.045 0.122 0.020 0.007 0.987 0.024 0.107 0.024 0.014 0.989 0.035 0.027 0.030 0.010 0.986 0.047 0.045
EDVR* [52] 0.036 0.014 0.979 0.038 0.119 0.015 0.007 0.989 0.020 0.106 0.009 0.010 0.988 0.033 0.027 0.028 0.009 0.986 0.049 0.043
VRT [53] 0.067 0.038 0.978 0.047 0.135 0.024 0.011 0.988 0.027 0.114 0.023 0.011 0.990 0.035 0.037 0.034 0.012 0.987 0.041 0.040
PhySR* [64] 0.084 0.036 0.966 0.040 0.113 0.051 0.022 0.979 0.037 0.099 0.032 0.011 0.988 0.029 0.026 0.041 0.014 0.984 0.045 0.038
Ours 0.028 0.011 0.981 0.033 0.100 0.014 0.006 0.990 0.011 0.093 0.007 0.002 0.996 0.014 0.016 0.018 0.005 0.991 0.035 0.028
@@
Visually, we report the super-resolution results of different methods in Figure 6 for the Bahamas dataset and in Figure 7 for the Galveston Bay. In Figure 6, the last row shows the ground truth. Denoting the simulation output for the n-th time step by $t_{n}$, we show the results for $t_{15}$ in the first column and the residual maps of $U$-, $V$- and $\xi$-planes between predictions and ground truth in columns 2, 3 and 4, respectively. The last column illustrates the residual maps between $t_{15}$ and $t_{30}$ to visualize the temporal changes. We can observe that 1) VDSR, RDN, EDVR, and PhySR wrongly predict the $V$-plane simulations, ST-Bicubic, SwinIR, and VRT wrongly predict the $U$-plane simulations, ResShift wrongly predicts both U- and $\xi$-planes. We can better minimize the simulation differences simultaneously for $U$-, $V$- and $\xi$-planes. 2) From the last column, we can see that ours can better match with the ground truth changes while others produce large errors. In Figure 7 of the super-resolution results on Galveston Bay (here, the ground truth is shown in the last column), we use white arrows to highlight the significant differences. We can see that 1) globally, ResShift and VRT enlarge the errors caused by ST-Bicubic (the green areas around the boundaries at time $t_{61}$). EDVR produces the wrong pattern at time $t_{41}$. 2) From the enlarged areas, we can see that ours can better mimic the changes of the solution around the corners. Others, like VDSR and PhySR* magnify the zigzag patterns. ResShift and VRT oversmooth the solution which visually causes color shifts.
@@
Figure 6: Visual comparison among different methods on Bahamas dataset. We show the ground truth at time $t_{15}$ and the corresponding residual maps between prediction and ground truth (U, V, and $\xi$ planes – we multiply the residuals by factor 50 to highlight the differences), and the residual map between the ground truths at times $t_{30}$ and $t_{15}$ (we multiply the residuals by factor 20 for visualization).
@@
Figure 7: Visual comparison among different methods on the Galveston dataset. The simulation results at time $t_{41}$ and $t_{61}$ in RGB images. We enlarge the red box region in rows 3 and 4 to highlight the visual differences.
@@
We report the number of parameters, FLOPs, and runtime for model complexity in Table II. The FLOPs are computed with the input of LR size 180 × 180 and ×4 upsampling settings. We can see that ours has the third lowest model complexity regarding the number of parameters and runtime. In general, the model performance and complexity contradict each other. To better visualize the trade-off comparisons among different methods, we summarize the information in Table III and Table II and visualize them in Figure 5. Ours is located at the bottom left corner, which indicates that it achieves the best balance between model complexity and reconstruction quality.
IV-C Ablation studies and analysis

### v2

@@ 4.2 Proposed DNNCS for coastal downscaling
In Figure 2, we show the overall structure of our proposed DNNCS . Given the input coarse simulation data $\mathbf{X}\in\mathbb{R}^{T\times H\times W\times C}$, where T is the number of adjacent views, H and W are the height and width, and $C=3$ is the $U$, $V$, $\xi$ planes. The proposed DNNCS learns the mapping function $f\colon\mathbf{X}\rightarrow\mathbf{Y}$, where $\mathbf{Y}\in\mathbb{R}^{\alpha T\times H\times W\times C}$ is the high-resolution coastal estimation, and $\alpha$ is the temporal upsampling factor. Note that we do not change the spatial resolution because the low-resolution and high-resolution coastal images are rendered from the graph with the same scale but with coarse and fine mesh resolutions, respectively. That is, the resolution of the simulation video does not change, but it becomes more detailed. Inside the network are three key components: Feature extraction, Spatiotemporal attention, and Feature split and reconstruction. Let us introduce them in detail. We train DNNCS on the Bahamas dataset at different downscaling scenarios, and then fine-tune it on the Galveston dataset. The detail of the proposed Bahamas and Galveston datasets will be introduced in Section 5.1
@@
Figure 2: The overall structure of the proposed DNNCS . We show the complete architecture of our proposed DNNCS. Given two consecutive coastal simulations, we take them as input to first extract the deep feature representation via multiple RCAB blocks. Then we use spatiotemporal attention to learn the pixel correlations across space and time. Finally, we split the features into three channels for temporal interpolation and spatial downscaling.
4.3 Feature extraction
@@ 4.4 Spatiotemporal attention
The key component of the DNNCS is the spatiotemporal attention. Our goal is to jointly super-resolve the coarse coastal images spatially and temporally. PDEs typically govern spatiotemporal dynamical systems. Let us first consider a simplified wave equation as
@@
4.5 Feature split and reconstruction
In the feature split and reconstruction process, we first project the learned features into the frequency domain using cosine and sine operators [^1] Note that we use cosine and sine operators for computational efficiency as they are used often in image/video coding to capture high-frequency details., then combine these components to enhance the original features. We subsequently apply a simple convolution layer to extend these features and split them into three subsets, each representing one of the three temporal feature maps for the coastal data. The input consists of two low-resolution coastal images, with the proposed network (DNNCS ) tasked with predicting the intermediate image. To achieve this, we use bilinear interpolation for temporal frame interpolation on the initial feature maps. A short connection is implemented to ensure that the spatiotemporal attention mechanism effectively learns the residuals. Finally, a shared convolution operation is employed to directly output the predicted downscaling coastal images. This approach ensures that the model captures both spatial and temporal details, producing high-quality intermediate frames.
4.6 Losses for coastal downscaling
To train the proposed DNNCS , we propose utilizing three loss terms to better constrain the visual consistency in space and time, including MAE (Mean Absolute Errors) loss, $L_{p}$ loss, and physics-informed loss. Given the ground truth and estimated downscaling results ${\mathbf{Y},\mathbf{Y^{\prime}}}\in\mathbb{R}^{\alpha T\times H\times W\times C}$, we first have the MAE loss as,
@@ 4.6 Losses for coastal downscaling
$\small L_{diff}=\left\|\frac{dY(t)}{dx}-\frac{dY^{\prime}(t)}{dx}\right\|^{2}+\left\|\frac{dY(t)}{dy}-\frac{dY^{\prime}(t)}{dy}\right\|^{2}+\left\|\frac{dY(t)}{dz}-\frac{dY^{\prime}(t)}{dz}\right\|^{2}$ (8)
@@
5 Experiments
5.1 Implementation Details
Datasets
To generate sufficient data samples for model training and analysis, we use UTBEST [1] to simulate the tidal circulation in two locations: Bahamas (Bight of Abaco) and Galveston Bay. The geometry of both computational domains, bathymetry (bottom topography), and the coarsest computational meshes are illustrated in Figure 4. Finer meshes were obtained by subdividing each triangle into four via edge bisection, whereas the bathymetry on finer meshes was obtained by linear interpolation from the coarser mesh nodes to preserve the consistency of problem setup.
@@ 5.1 Implementation Details
At the open sea boundary, both test problems use the tidal elevation consisting of five constituents prescribed by the function
@@
$+0.1\cos\left(\frac{t}{12.66}+5.93\right)+0.395\cos\left(\frac{t}{12.42}\right)+0.06\cos\left(\frac{t}{12.00}+0.75\right)$
with time $t$ in hours and use a quadratic bottom friction $\tau_{bf}=C_{f}|\mathbf{q}|/H^{2}$ with $C_{f}=0.009$ (Bahamas) and $C_{f}=0.004$ (Galveston). The Coriolis parameter for Bahamas and Galveston is set to $3.19\times 10^{-5}~{}\mathrm{s}^{-1}$ and $7.07\times 10^{-5}~{}\mathrm{s}^{-1}$, respectively (see [1] for a more detailed description of the test problems).
The simulations are run for 24 hours with the time step for the lowest approximation order (piecewise constant DG spaces) and the lowest mesh resolution set to 20s for the Bahamas and 12s for the Galveston dataset. Output files with the water height $\xi$ and depth-integrated horizontal velocities $U$ and $V$ are generated in each time step for all configurations. Increasing the mesh resolution or the polynomial discretization order requires smaller time steps (we halve the time step for each mesh refinement and each order increment) and results in more output files. To provide paired LR and HR training data, we consider three levels of downscaling scenarios: size of simulation data, polynomial order of mesh reconstruction and the temporal sampling. Table 1 summarizes the data used for training, with the number of mesh elements shown in “Resolution” and the number of time steps in the “Order” column. We use three approximation orders and resolutions to produce corresponding coarse and fine simulation data at different temporal resolutions. That is, given three approximation orders and three different resolutions, we have a total of $3\times 3$ simulation cases for each location. Then, we render all data of their three views ($U$, $V$, $\xi$ planes) into grayscale images. Finally, all three views are merged as RGB images as the training data. We train one model on both the Bahamas and Galveston datasets for our evaluation.
Table 1: Summary of the data used in coastal simulation for Bahamas and Galveston. The “Resolution” value represents the number of mesh elements. In the “Order” column, we list the number of time steps in the 24-hour simulation for the corresponding polynomial discretization order.
@@
We randomly split the data into training, validation, and testing in $6:2:2$ ratios without overlapping. The rendered images are of identical dimensions, approximately $900\times 500$ pixels. We also mask out the land area and only focus on the water-covered subdomain downscaling. In the training phase, we learn 4$\times$ spatial downscaling and 2$\times$ temporal downscaling across different orders. We randomly select two adjacent LR coastal images ($\mathbf{X}(t_{n}),\mathbf{X}(t_{n+1})$) from lower resolutions to predict three corresponding HR images $\mathbf{Y}(t_{2n}),\mathbf{Y}(t_{2n+1}),\mathbf{Y}(t_{2n+2})$ from next higher-resolution. Then we crop the image pairs as patches of size $64\times 64$. We randomly apply horizontal & vertical flips, rotations, and frame reversal for data augmentation.
Parameter setting
@@
We train DNNCS using Adam optimizer with the learning rate of $1\times 10^{-4}$. The learning rate is halved after 30k iterations. The batch size is set to 24, and DNNCS is trained for 100k iterations (about 16 hours) on a PC with one NVIDIA V100 GPU using the PyTorch deep learning platform.
Metrics and evaluation
We evaluate different models separately for two scenarios: Bahamas and Galveston. In each scenario, we average the model predictions at different orders. We use five metrics for evaluation:
  • 1.
    MSE (Mean Squared Error) and MAE (Mean Absolute Error) measure the average pixel differences between ground truth and estimation
  • 2.
    SSIM [78] measures the structural similarity between ground truth and estimation
  • 3.
    GMSD [81] measures the gradient differences for perceptual quality assessment
  • 4.
    LPIPS [83] measures the perceptual similarity between ground truth and estimation
To measure the model complexity, we use FLOPs, memory, and the number of parameters to compare the computation efficiency.
5.2 Comparison of the spatiotemporal downscaling with state-of-the-art methods
@@
To show the efficiency of our proposed method, we compare it with four state-of-the-art image super-resolution methods: VDSR [46], RDN [85], StableSR [75] and SwinIR [50], two video super-resolution methods: EDVR [76] and VRT [49], and one physics-informed downscaling method: PhySR [59]. Note that image super-resolution cannot conduct frame interpolation, so we apply spatiotemporal Bicubic to interpolate the missing frames and then apply frame-by-frame downscaling. All approaches are reimplemented using the publicly available codes, and the methods marked with * indicate that they were retrained from our coastal dataset for a fair comparison. In Table 2, both Galveston and Bahamas have two downscaling schemes. The “Resolution” column lists the number of grid elements. The result for each scheme is the average of all three polynomial approximation orders. We can see that our approach achieves the best performance with respect to the RMSE, MAE, SSIM, and GMSD metrics and is the second-best in LPIPS.
@@
Table 2: Comparison with state-of-the-art methods on coastal downscaling. We report the deviations from the ground truth given by the highest accuracy simulations for the test datasets of Bahamas and Galveston using different metrics.
Dataset Resolution Method ST-Bicubic VDSR* RDN* ResShift SwinIR EDVR* VRT PhySR* Ours
@@
[…]
@@
Visually, we report the downscaling results of different methods in Figure 5 for the Bahamas dataset and in Figure 6 for the Galveston Bay. Each plane is normalized between [0, 255] and shown as a single-channel grayscale image, and we also stack normalized three planes $U,V,\xi$ in order as RGB images. In Figure 5, the last row shows the ground truth. Denoting the simulation output for the n-th time step by $t_{n}$, we show the results for $t_{15}$ in the first column and the residual maps of $U$-, $V$- and $\xi$-planes between predictions and ground truth in columns 2, 3 and 4, respectively. The last column illustrates the residual maps between $t_{15}$ and $t_{30}$ to visualize the temporal changes. We can observe that 1) VDSR, RDN, EDVR, and PhySR wrongly predict the $V$-plane simulations, ST-Bicubic, SwinIR, and VRT mispredict the $U$-plane simulations, ResShift fails on both U- and $\xi$-planes. We can better minimize the simulation differences simultaneously for $U$-, $V$- and $\xi$-planes. 2) From the last column, we can see that ours can better match with the ground truth changes, while others produce large errors. In Figure 6 of the downscaling results on Galveston Bay (here, the ground truth is shown in the last column), we use white arrows to highlight the significant differences. We can see that 1) globally, ResShift and VRT enlarge the errors caused by ST-Bicubic (the green areas around the boundaries at time $t_{61}$). EDVR produces the wrong pattern at time $t_{41}$. 2) From the enlarged areas, we can see that ours can better mimic the changes of the solution around the corners. Others, like VDSR and PhySR* magnify the zigzag patterns. ResShift and VRT oversmooth the solution which visually causes color shifts.
Three main reasons that other methods fail while ours work: 1) Other image-based approaches (SwinIR, ResShift) are trained with additional perceptual reconstruction losses, which can cause false features or hallucinate detail. 2) Video-based approaches (EDVR, VRT) use estimated optical flow generated by pre-trained neural networks for temporal interpolation. The errors caused by optical flow can be accumulated for spatial downscaling. 3) All other works do not consider physics-informed losses to ensure physical consistency, which would violate the energy flow between frames.
@@
Figure 5: Visual comparison among different methods on Bahamas dataset. We show the ground truth from resolution 1696, order 0 at time $t_{15}$ and $t_{30}$, and the corresponding residual maps between prediction and ground truth (U, V, and $\xi$ planes – we multiply the residuals by factor 50 to highlight the differences), and the residual map between the ground truths at times $t_{30}$ and $t_{15}$ (we multiply the residuals by factor 20 for visualization).
@@
Figure 6: Visual comparison among different methods on the Galveston dataset. We use the data sample from resolution 3397, order 1 at time $t_{41}$ and $t_{61}$. The simulation results at time $t_{41}$ and $t_{61}$ in RGB images. We enlarge the red box region in rows 3 and 4 to highlight the visual differences.
@@
We report the number of parameters, FLOPs, and runtime for model complexity in Table 3. The FLOPs are computed with the input of LR size 180 × 180 and ×4 upsampling settings. We can see that ours has the third-lowest model complexity regarding the number of parameters and runtime. In general, the model performance and complexity contradict each other. To better visualize the trade-off comparisons among different methods, we summarize the information in Table 2 and Table 3 and visualize them in Figure 7. Ours is located at the bottom left corner, which indicates that it achieves the best balance between model complexity and reconstruction quality.
Table 3: Comparison with state-of-the-art methods on model complexity. We report the results on the Bahamas dataset under the same hardware settings.
Method ST-Bicubic VDSR* RDN* ResShift SwinIR EDVR* VRT PhySR* Ours
Number of parameters (M) - 0.651 21.99 118.59 11.72 20.6 35.5 0.817 8.07
FLOPs (GMac) - 8.23 90.14 32156 1878.41 3020.1 3305.75 9.20 65.15
Runtime (s) 0.003 0.038 0.48 15.2 4.2 1.10 2.4 0.23 1.02
Figure 7: Speed and performance comparison. Overall comparison among different methods in terms of speed (runtime on the vertical axis), performance (RMSE on the horizontal axis), and model complexity (size of the bubbles).
5.3 Ablation studies and analysis
Losses for optimization